# SDTrack: A Baseline for Event-based Tracking via Spiking Neural Networks

### v1

@@ 3 SDTrack Pipeline
3.1 I-LIF Spiking Neuron
Recently, the emergence of I-LIF neurons enhances the expressiveness of SNNs [23, 34, 20, 25, 39]. This enhancement enables SNNs to potentially outperform ANNs in various machie learning tasks [23, 20, 34], while preserving their spike-driven characteristics. Therefore, we employ the I-LIF spiking neuron model to build our SDTrack pipeline. I-LIF neurons transmit information through spike discharge frequency in the training stage, while utilizing binary spike signals during the inference stage [23, 39]. In the training stage, the output of the I-LIF neuron can be expressed as
${\mathbf{s}^{\ell}}=\frac{1}{T}\cdot\left\lfloor\operatorname{clip}\left\{{\mathbf{x}^{\ell}},\text{0},T\right\}\right\rceil,$ (1)
[…]
In the inference stage, the firing rate $\mathbf{s}^{\ell}$ is decoupled into $T$ binary spike signals, mathematically described as
${\mathbf{s}^{\ell}}=\frac{1}{T}\cdot\sum\limits_{t=1}^{T}{\mathbf{s}^{\ell}[t]}.$ (2)
At this stage, for neurons in layer $l+1$, whether considering matrix multiplication or Hadamard product as linear operators, their input can be expressed as
${\mathbf{x}^{\ell+1}}={\mathbf{w}^{\ell}}\cdot{\mathbf{s}^{\ell}}=\sum\limits_{t=1}^{T}{\frac{{{\mathbf{w}^{\ell}}}}{T}}{\mathbf{s}^{\ell}}[t],$ (3)
where $\mathbf{w}^{\ell}$ is the weight matrix in linear or convolutional layer, ${\mathbf{s}^{\ell}}[t]\in\{0,1\}$ is a binary spike at time $t$. Therefore, I-LIF maintains the key spike-driven property of SNNs during inference, which enables the transformation of MAC into efficient AC operations, significantly reducing storage and computational overhead during edge deployment.
@@ 3.2 Global Trajectory Prompt
where $L$ is the duration of each event stream and $\alpha$ is a coefficient that preserves more valid data while removing partial noise. By using our GTP, the first and second channels in a event image, i.e., $h_{i}^{1}$ and $h_{i}^{2}$, accumulate the number of positive and negative polarities within the event stream, respectively. As shown in Fig 3, when multiple directional movements occur at a pixel within the interval $L$, prior methods only select the most recent polarity value as the aggregation result, discarding preceding motion information. In contrast, GTP separately accumulates the number of positive and negative polarities, thus preserving all motion information throughout the entire time interval.
@@ 3.3 Transformer-based Spike-driven Tracker
where $\mathbf{U}\in(T,C,{H_{z}+H_{x}},{W_{z}+W_{x}})$. As demonstrated in Tab. 2, our approach enables SDTrack to capture stronger positional information compared to conventional encoding schemes, without explicitly requiring positional encoding. Due to the spike-driven nature of our architecture, the additional padding introduces negligible computational overhead.
@@
Tracking Head. Similar to many trackers, SDTrack uses a simple center tracking head to directly predict the target position and size. What distinguishes SDTrack is its SNN-redesigned center head. In each branch, the first four convolutional operations are based on spike signals while the final convolutional operation is based on floating-point signals. This design addresses the inherent limitation of spike-based convolutions, which are insufficient for precise object localization and scale estimation—a deficiency that cannot be compensated through continuous learning. We demonstrate the necessity of this approach for constructing SNN-based trackers through both experimental validation in Sec 5 and theoretical derivation in the Supplementary Materials. Furthermore, in Sec 4.4, we demonstrate that center heads are more suitable for SDTrack compared to corner heads.
@@
where ${\lambda_{iou}}=2$ and ${\lambda_{L1}}=5$ are the regularization parameters in our experiments, following [35]. The detailed description of the loss function is provided in the Supplementary Material. Since neurons transmit spike firing frequencies during the training stage, the network’s simulation time step is set to 1.
During inference, we follow the standard SOT procedure by using the first frame of the test sequence as the template. Notably, we do not employ dynamic template strategies or post-processing techniques such as hanning window penalty. Our GTP method is applicable to real-world scenarios in real-time, as illustrated by the purple box in Fig. 2. Notably, during inference, the network’s simulation time step is set to T, as the spike firing rate used in training is systematically decoupled into precise spike signals over T time steps according to Eq. 2.
Methods Parameters (M) Timestep Power (mJ) FE108 [43] FELT [29] VisEvent [28]
@@
STARK [35] 28.23 1 58.88 57.4 89.2 39.6 51.7 34.1 46.8
SimTrack [4] 88.64 1 93.84 56.7 88.3 36.8 47.0 34.6 47.6
OSTrack256 [40] 92.52 1 98.90 54.6 87.1 35.9 45.5 32.7 46.4
ARTrack256 [32] 202.56 1 174.80 56.6 88.5 39.5 49.4 33.0 43.8
SeqTrack-B256 [6] 90.60 1 302.68 53.5 85.5 33.0 42.0 28.6 43.3
HiT-B [17] 42.22 1 19.78 55.9 88.5 38.5 48.9 34.6 47.6
GRM [12] 99.83 1 142.14 56.8 89.3 37.2 47.4 33.4 47.7
HIPTrack [3] 120.41 1 307.74 50.8 81.0 38.2 48.9 32.1 45.2
ODTrack [47] 92.83 1 335.80 43.2 69.7 29.7 35.9 24.7 34.7
SiamRPN* [21] - 1 - - - - - 24.7 38.4
ATOM* [9] - 1 - - - 22.3 28.4 28.6 47.4
DiMP* [2] - 1 - - - 37.8 48.5 31.5 44.2
PrDiMP* [10] - 1 - - - 34.9 44.5 32.2 46.9
MixFormer* [8] 37.55 1 - - - 38.9 50.4 - -
STNet* [44] 20.55 3 - - - - - 35.0 50.3
SNNTrack* [46] 31.40 5 8.25 - - - - 35.4 50.4
SDTrack-Tiny 19.61 4 8.16 59.0 91.3 39.6 50.1 35.6 49.2
SDTrack-Base 107.26 4 30.52 59.9 91.5 39.8 50.7 37.4 51.5
Table 1: Comparison with standard SOT pipeline on three event-based tracking benchmarks. Due to incompatibility with the GTP methodology, ODTrack utilizes the data processing approach proposed by Wang et al. for comparative analysis [28]. Works marked with * indicate direct adoption of results from their respective benchmark reports. The remaining results employ the GTP methodology for event aggregation to better demonstrate SDTrack’s performance advantages. The energy consumption of SDTrack is represented by its measurements on the VisEvent dataset. The best two results are shown in red and blue fonts.
@@ 4.2 Evaluation Metric
$+{E_{\operatorname{AC}}}\cdot(\sum\limits_{j}{\operatorname{FL}_{\operatorname{SNNConv}}^{j}\cdot f{r^{j}}+\sum\limits_{k}{\operatorname{FL}_{\operatorname{SNNFC}}^{k}\cdot f{r^{k}}}})),$
where $i$ represents every non spike-driven convolution layer and $j,k$ represent every spike-driven convolutional and fully-connected (FC) layers, respectively. Let ${E_{\operatorname{MAC}}}$ and ${E_{\operatorname{AC}}}$ denote the energy costs of MAC and AC operations, while ${FL_{\operatorname{SNNConv}}^{j}}$ and ${FL_{\operatorname{SNNFC}}^{k}}$ are the FLOPs of $j$-th convolutional layer and $k$-th FC layer. Additionally, ${f{r^{j}}}$ and ${f{r^{k}}}$ denote the firing rate (the proportion of non-zero elements) of the corresponding spike matrices at matching positions. Following the established methodology for calculating energy consumption in prior SNNs studies [38, 37, 39], we assume the data for various operations are 32-bit floating-point implementation in 45nm technology [15], in which ${E_{\operatorname{MAC}}}=4.6pJ$ and ${E_{\operatorname{AC}}}=0.9pJ$. The detailed computation methods for all metrics and the firing rate of SDTrack are presented in the Supplementary Materials.
@@ 4.3 Comparison with Standard Pipeline
Event-based Tracker. Building upon these findings, we further evaluate the effectiveness of the SDTrack architecture on event-based tracking datasets FE108 [43], VisEvent [28] and FELT [29], with detailed dataset specifications provided in the Supplementary Materials. Notably, to ensure fair comparison without the overwhelming advantages of GTP methods, most baseline approaches also utilize GTP for event aggregation, thus demonstrating SDTrack’s inherent advantages in event-based tracking tasks.
@@
As shown in Tab. 1 and Fig. 5, on all three datasets, SDTrack demonstrates SOTA potential. On the FE108 dataset, SDTrack-Tiny achieves significant improvements over the previous best model GRM, with a 4.0% increase in AUC and a 2.2% gain in PR. The Base model further enhances these performance metrics compared to its Tiny counterpart. On the FELT dataset, SDTrack outperforms STARK in AUC metrics while achieving the second-highest score in PR metrics. Notably, STARK’s PR performance relies on the GTP method; without it, STARK’s PR score drops to 50.8, comparable to SDTrack-Base. On the crucial VisEvent dataset, which presents higher complexity, SDTrack-Tiny achieves PR scores slightly below STNet and SNNTrack. However, its AUC performance remains comparable to the previous best model, SNNTrack. The larger-scale SDTrack-Base model surpasses SNNTrack’s previous state-of-the-art results, demonstrating a 2.0% improvement in AUC and a 1.1% increase in PR.
Notably, while SDTrack’s parameter count approximates that of previous Base-level models, its theoretical energy consumption on edge devices is substantially lower than models of comparable scale. Compared to HiT-B, which was specifically designed for lightweight deployment, our SDTrack-Tiny significantly outperforms across multiple datasets while utilizing less than half the parameters and energy consumption. These results substantiate both the energy efficiency characteristics of SNNs and their potential for edge deployment. Furthermore, due to the natural compatibility between SNNs and event data, coupled with the efficient design of our SDTrack pipeline, SNNs demonstrate the potential to surpass ANNs in event tracking tasks.
Figure 6: Selection of GTP hyperparameters. The entire selection process is conducted using SDTrack-Tiny on the FE108 dataset. (a) shows the selection experiment for $\alpha$, where events are aggregated into a two-channel image. (b) illustrates the selection process for $\beta$, during which the $\alpha$ in GTP is consistently set to 30.
@@ 4.4 Ablation Study
GTP. The GTP method involves two hyperparameters: $\alpha$ and $\beta$. Through experimental evaluation on the FE108 dataset, we determined that setting $\alpha$ to 30 and $\beta$ to 0.8 is the optimal configuration. The experimental process is illustrated in Fig 6.
@@
Table 2: Ablation studies on the FE108 dataset.

### v2

@@ 3 SDTrack Pipeline
3.1 Unified Spiking Neuron Model
Fundamental dynamics of postsynaptic neurons include membrane potential elevation after receiving presynaptic neuron spikes, spike generation upon membrane potential reaching threshold, and membrane potential reset following spike emission. Since our method is general and independent of specific neuron types, we present a unified dynamical equation for spiking neurons, described as
$\mathbf{U}[t]$ $=\mathbf{H}[t-1]+\frac{1}{\tau}(\mathbf{X}[t]-(\mathbf{H}[t-1]-{\mathbf{U}_{reset}})),$ (1)
$\mathbf{S}[t]$ $=f(\mathbf{U}[t]-{\mathbf{U}_{threshold}}),$ (2)
$\mathbf{H}[t]$ $=\mathbf{U}[t](1-\mathbf{S}[t]).$ (3)
For any neuron, $\mathbf{U}[t]$ represents its membrane potential at time $t$ after receiving input $\mathbf{X}[t]$, while $\mathbf{S}[t]$ denotes the neuron’s output at time $t$, and $\mathbf{H}[t]$ signifies the residual membrane potential following the completion of the neuron’s lifecycle at time $t$. The parameter $\tau$ functions as the decay factor of the neuronal membrane potential, $\mathbf{U}_{reset}$ represents the reset membrane potential post-firing, and $\mathbf{U}_{threshold}$ denotes the firing threshold of the neuron. These three parameters constitute the adjustable hyperparameters within the model architecture.
Both IF (Integrate-and-Fire) [3] and LIF (Leaky Integrate-and-Fire) [23] neurons, the discrimination function $f(\cdot)$ is implemented as a Heaviside step function. In this formulation, when $\tau$ equals 1, the neuron functions as an IF neuron; when $\tau$ exceeds 1, it operates as a LIF neuron.
This model represents an I-LIF neuron [37] when parameters $\mathbf{U}_{reset}$ and $\mathbf{U}_{threshold}$ are both set to 0 and $f(x)$ corresponds to ${\frac{1}{D}\cdot\operatorname{Clip}(\operatorname{round}(x),0,D)}$, where $x$ always represents $\mathbf{U}[t]$, $D$ denotes virtual time steps, $\operatorname{clip}(x,min,max)$ constrains $x$ to the range $[min,max]$, and $\operatorname{round}(\cdot)$ performs the rounding operation. During inference, this model can be converted to 0/1 spike signals using the spike-ahead principle, where $D$ is incorporated into actual iterative timestep $T$, forming a total of $T\times D$ time steps. Therefore, all these neurons achieve spike-driven implementation during the inference phase, and neurons with this characteristic replace MAC operations with simpler AC operations during inference, significantly reducing storage and computational costs for edge deployment.
@@ 3.2 Global Trajectory Prompt
where $L$ is the duration of each event stream and $\alpha$ is a coefficient that preserves more valid data while removing partial noise. By using our GTP, the first and second channels in a event image, i.e., $h_{i}^{1}$ and $h_{i}^{2}$, accumulate the number of positive and negative polarities within the event stream, respectively. As shown in Fig. 3, when multiple directional movements occur at a pixel within the interval $L$, prior methods only select the most recent polarity value as the aggregation result, discarding preceding motion information. In contrast, GTP separately accumulates the number of positive and negative polarities, thus preserving all motion information throughout the entire time interval.
@@ 3.3 Transformer-based Spike-driven Tracker
[…]
where $\mathbf{U}\in(T,C,{H_{z}+H_{x}},{W_{z}+W_{x}})$, ${{\rm O}_{1}}$ and ${{\rm O}_{2}}$ represent zero tensors. As demonstrated in Tab. 2, our approach enables SDTrack to capture stronger positional information compared to conventional encoding schemes, without explicitly requiring positional encoding. Due to the spike-driven nature of our architecture, the additional padding introduces negligible computational overhead.
@@
Tracking Head. Similar to many trackers, SDTrack uses a simple center tracking head to directly predict the target position and size. What distinguishes SDTrack is its SNN-redesigned center head. In each branch, the first four convolutional operations are based on spike signals while the final convolutional operation is based on floating-point signals. This design addresses the inherent limitation of spike-based convolutions, which are insufficient for precise object localization and scale estimation—a deficiency that cannot be compensated through continuous learning. We demonstrate the necessity of this approach for constructing SNN-based trackers through both experimental validation in Sec. 5 and theoretical derivation in the Supplementary Materials. Furthermore, in Sec. 4.4, we demonstrate that center heads are more suitable for SDTrack compared to corner heads.
@@
where ${\lambda_{iou}}=2$ and ${\lambda_{L1}}=5$ are the regularization parameters in our experiments, following [33]. The detailed description of the loss function is provided in the Supplementary Material.
During inference, we follow the standard SOT procedure by using the first frame of the test sequence as the template. Notably, we do not employ dynamic template strategies or post-processing techniques such as hanning window penalty. Our GTP method is applicable to real-world scenarios in real-time, as illustrated by the purple box in Fig. 2.
Methods Param. (M) Spiking Neuron Timesteps ($T\times D$) Power (mJ) FE108 [41] FELT [28] VisEvent [27]
@@
STARK [33] 28.23 - $1\times 1$ 58.88 57.4 89.2 39.3* 50.8* 34.1 46.8
SimTrack [5] 88.64 - $1\times 1$ 93.84 56.7 88.3 36.8 47.0 34.6 47.6
OSTrack256 [38] 92.52 - $1\times 1$ 98.90 54.6 87.1 35.9 45.5 32.7 46.4
ARTrack256 [31] 202.56 - $1\times 1$ 174.80 56.6 88.5 39.5 49.4 33.0 43.8
SeqTrack-B256 [7] 90.60 - $1\times 1$ 302.68 53.5 85.5 33.0 42.0 28.6 43.3
HiT-B [18] 42.22 - $1\times 1$ 19.78 55.9 88.5 38.5 48.9 34.6 47.6
GRM [13] 99.83 - $1\times 1$ 142.14 56.8 89.3 37.2 47.4 33.4 47.7
HIPTrack [4] 120.41 - $1\times 1$ 307.74 50.8 81.0 38.2 48.9 32.1 45.2
ODTrack [45] 92.83 - $1\times 1$ 335.80 43.2 69.7 29.7 35.9 24.7 34.7
SiamRPN* [21] - - $1\times 1$ 24.7 38.4
ATOM* [10] - - $1\times 1$ 22.3 28.4 28.6 47.4
DiMP* [2] - - $1\times 1$ 37.8 48.5 31.5 44.2
PrDiMP* [11] - - $1\times 1$ 34.9 44.5 32.2 46.9
MixFormer* [9] 37.55 - $1\times 1$ 38.9 50.4 - -
STNet* [42] 20.55 LIF $3\times 1$ 35.0 50.3
SNNTrack* [44] 31.40 BA-LIF $5\times 1$ 8.25 - - - - 35.4 50.4
SDTrack-Tiny 19.61 LIF $4\times 1$ 8.15 56.7 89.1 35.8 44.0 35.4 48.7
I-LIF $2\times 2$ 9.87 55.3 88.1 35.7 45.3 35.4 49.5
I-LIF $1\times 4$ 8.16 59.0 91.3 39.3 51.2 35.6 49.2
SDTrack-Base 107.26 I-LIF $1\times 4$ 30.52 59.9 91.5 40.0 51.4 37.4 51.5
Table 1: Comparison with standard SOT pipeline on three event-based tracking benchmarks. Due to incompatibility with the GTP methodology, ODTrack utilizes the data processing approach proposed by Wang et al. for comparative analysis [27]. * denotes results directly adopted from their respective benchmark reports. The remaining results employ the GTP methodology for event aggregation to better demonstrate SDTrack’s performance advantages. The energy consumption of SDTrack is represented by its measurements on the VisEvent dataset. The best two results are shown in red and blue fonts.
@@ 4.2 Evaluation Metric
$+{E_{\operatorname{AC}}}\cdot(\sum\limits_{j}{\operatorname{FL}_{\operatorname{SNNConv\&FC}}^{j}\cdot f{r^{j}})+\sum\limits_{k}{\operatorname{FL}_{\operatorname{SSA}}^{k}}}),$
where ${E_{\operatorname{MAC}}}$ and ${E_{\operatorname{AC}}}$ denote the energy costs of MAC and AC operations, respectively, while $\operatorname{FL}_{\operatorname{SNNConv\&FC}}$ represents the FLOPs of the $j$-th convolutional layer or FC layer, and $\operatorname{FL}_{\operatorname{SSA}}^{k}$ represents the FLOPs of the $k$-th SNN Self-Attention module. Additionally, ${f{r^{j}}}$ denote the firing rate (the proportion of non-zero elements) of the corresponding spike matrices at matching positions. Following the established methodology for calculating energy consumption in prior SNNs studies [36, 35, 37], we assume the data for various operations is implemented using 32-bit floating-point precision in 45nm technology [16], where ${E_{\operatorname{MAC}}}=4.6pJ$ and ${E_{\operatorname{AC}}}=0.9pJ$. The detailed computation methods for all metrics and the firing rate of SDTrack are presented in the Supplementary Materials.
@@ 4.3 Comparison with Standard Pipeline
Event-based Tracker. Based on these findings, we further evaluate the effectiveness of SDTrack models of different sizes equipped with different neurons on event-based tracking datasets FE108 [41], VisEvent [27], and FELT [28], and provide detailed dataset specifications in the Supplementary Material. Notably, even though most baseline methods utilize GTP methods with significant performance advantages for event aggregation, SDTrack models still outperform them.
@@
As shown in Tab. 1 and Fig. 5, across all three datasets, the SDTrack-Tiny model equipped with I-LIF neurons achieves the lowest parameter count and energy consumption. On the FE108 dataset, it improves AUC by 1.6% and PR by 2.0% compared to the previous best model, and demonstrates SOTA potential on the FELT and VisEvent datasets. The Tiny models equipped with LIF and temporally-featured I-LIF neurons ($T=2\times 2$) show SOTA potential on the FE108 and VisEvent datasets, but perform poorly on the long sequence tracking benchmark FELT dataset. Compared to the Tiny models, the Base models further improve these performance metrics, achieving SOTA results on both AUC and PR metrics across all three event-based tracking datasets.
Notably, while SDTrack-Base maintains a parameter count comparable to previous Base-level models, it achieves substantially lower theoretical energy consumption on edge devices. Compared to HiT-B, which was specifically designed for lightweight deployment, our SDTrack-Tiny significantly outperforms across multiple datasets while utilizing less than half the parameters and energy consumption. These results substantiate both the energy efficiency characteristics of SNNs and their potential for edge deployment. Furthermore, due to the natural compatibility between SNNs and event data, coupled with the efficient design of our SDTrack pipeline, SNNs demonstrate the potential to surpass ANNs in event tracking tasks.
Figure 6: Selection of GTP hyperparameters. (a) shows the selection experiment for $\alpha$, where events are aggregated into a two-channel image. (b) illustrates the selection process for $\beta$, during which the $\alpha$ in GTP is consistently set to 30.
@@ 4.4 Ablation Study
All ablation experiments follow the same experimental setup and are conducted using SDTrack-Tiny equipped with I-LIF neurons of $T=1\times 4$ on FE108.
GTP. The GTP method involves two hyperparameters: $\alpha$ and $\beta$. Through experimental evaluation on the FE108 dataset, we determined that setting $\alpha$ to 30 and $\beta$ to 0.8 is the optimal configuration. The experimental process is illustrated in Fig. 6.
@@
Table 2: Ablation studies.